OpenStudy (anonymous):

Use the formula to evaluate the infinite series. 25+5+1+...

OpenStudy (anonymous):

the formula is

OpenStudy (anonymous):



hartnn (hartnn):

a1 = 1st term = 25 can you find the common ratio =r =... ?

OpenStudy (anonymous):

would it be 5

hartnn (hartnn):

the common ratio is next term/ current term = 5/25 = ... ?

OpenStudy (anonymous):

5/25=0.20

hartnn (hartnn):

yeah, use that

OpenStudy (anonymous):



OpenStudy (anonymous):

So then the answer would be 31.25